OpenStudy (nuccioreggie):

Line H is represented by the following equation: 2x + 2y = 8 What is most likely the equation for line K so the set of equations has infinitely many solutions? x + 2y = 2 x + y = 4 2x + y = 6 x − y = 8

OpenStudy (nuccioreggie):

i think its B

OpenStudy (whovianchick):

To do this one, you need to figure out which equation is the same equation as the given. Given equation: 2x + 2y = 8 Well, you can see off the bat that B is going to be the answer. x+y=4 2(x)=2(y)=2(4) 2x+2y=8 see? but it's good to check the others if it's not so obvious. Do that by lining them up. 2x + 2y = 8 x + 2y = 2 You can see it won't be A, because the x would have to multiplied by 2, while the y would stay the same, and the plain number on the other side of the equal sign would have to be multiplied by 4. The most important rule with equations is this: whatever you do to one side of the equation, you have to do to both. You could only multiply the equation by one thing, and you would have to do it to every single variable in the equation. 2x+2y=8 2x + y = 6 Again, same thing. the x would have to stay the same, the y and 6 would have to be multiplied by different things, and that wouldn't work. You know that D is wrong because there's a negative on the left side of the equation without a negative added to the right side of the equation.

OpenStudy (whovianchick):

Yo what level of math are you? This sounds a bit like Algebra I

OpenStudy (nuccioreggie):

so its C ?

OpenStudy (nuccioreggie):

and 8th

OpenStudy (whovianchick):

no it's B, just like I said.

OpenStudy (whovianchick):

Are you reading the responses I'm giving? I'm trying to help you understand why I'm doing what I'm doing.

OpenStudy (nuccioreggie):

oh sorry i ment to put B not c lol my bad

OpenStudy (whovianchick):

lol ur good I'll be right back I have to take tylenol my braces hurt

OpenStudy (nuccioreggie):

ok